an ice cream cone of radius 3cm hold 100ml of ice cream. what is the height of the cone
Answer by Boreal(15235)

v of cone=(1/3)pi*r^2*h
100=(1/3)pi^3^2*h
300/9*pi=h
h=10.61 cm